OpenStudy (anonymous):

Elena gives the same amount of water to each of her 7 rose bushes. She has 5 gallons of water for the bushes. How much water does Elena give to each rose bush? __ gal

OpenStudy (anaise):

What grade?

OpenStudy (anaise):

This is pretty easy :)

OpenStudy (anonymous):

5th

OpenStudy (anonymous):

sisters and I did not understand

OpenStudy (anaise):

Makes sense, can you divide 5 by 7?

OpenStudy (anonymous):

It has to be in a fraction

OpenStudy (anaise):

Yup

OpenStudy (anonymous):

5divided by 7 is 0.71428571428571

OpenStudy (anaise):

XD

OpenStudy (anonymous):

confused

OpenStudy (anaise):

The calculator you are using will keep rounding it off.

OpenStudy (anonymous):

oh

OpenStudy (kendricklamar2014):

Divide the amount of rose bushes (7) by how many gallons of water she has (5). \[\frac{ 7 }{ 5 } = \] @DAVPAS1

OpenStudy (anaise):

Ken, leave.

OpenStudy (kendricklamar2014):

?

OpenStudy (anaise):

Get away from this post

rebeccaxhawaii (rebeccaxhawaii):

no i want Ken here

OpenStudy (anaise):

I did not ask you

OpenStudy (anaise):

Yuck

OpenStudy (kendricklamar2014):

@DAVPAS1 What is: \[\frac{ 7 }{ 5 }\]

rebeccaxhawaii (rebeccaxhawaii):

i just telling you now so you dont ask later

OpenStudy (anaise):

Ken, she probably needs the explanation and guidance :)

OpenStudy (anaise):

They're using google calx

OpenStudy (kendricklamar2014):

Thats what I did....

OpenStudy (anaise):

Calc* and no

OpenStudy (kendricklamar2014):

Yes I am. This is what I said: Divide the amount of rose bushes (7) by how many gallons of water she has (5).

OpenStudy (anaise):

I am already peeved *rage quits* Your answer is 1 3/4

OpenStudy (anaise):

Oh

OpenStudy (anaise):

Nvm, I thought it was a 4

OpenStudy (anaise):

It would be 1 and 2/5 Tell me I'm wrong.

OpenStudy (anonymous):

thx @KendrickLamar2014

OpenStudy (kendricklamar2014):

It would, but u need to simplify the answer

OpenStudy (anaise):

What the hell

OpenStudy (kendricklamar2014):

\[1 \frac{ 2 }{ 5 } = \frac{ 7 }{ 5 } = 1.4\]

OpenStudy (kendricklamar2014):

The answer would be 1.4 gallons

OpenStudy (kendricklamar2014):

Welcome =)

OpenStudy (anaise):

She needed it in that form? Omg

OpenStudy (anaise):

Ken, medal me.

OpenStudy (anonymous):

ok

OpenStudy (kendricklamar2014):

@DAVPAS1 does the answer have to be a fraction or a decimal?

OpenStudy (anaise):

Good good

OpenStudy (anonymous):

fraction

OpenStudy (anaise):

I WAS RIGHT

OpenStudy (kendricklamar2014):

Ok, so the answer would be: \[\frac{ 7 }{ 5 }\]

OpenStudy (anaise):

1 2/5

OpenStudy (anonymous):

wait what 7/5 XD

OpenStudy (anonymous):

or 1 2/5

OpenStudy (anaise):

Simplified version, you gave ken the medal, which is wrong

OpenStudy (anonymous):

wait its the same

OpenStudy (kendricklamar2014):

How did you get \[1 \frac{ 2 }{ 5 }\] @Anaise

OpenStudy (anonymous):

7/5 simplified

OpenStudy (anaise):

I did exactly what you did, just

OpenStudy (kendricklamar2014):

Yes they are the same, but I want to know what she did to get that.

OpenStudy (anonymous):

well to get 1 2/5 it is = 7/5

OpenStudy (anaise):

I learn a different way excuse me, but I subtracted 5 from 7, made that the whole and put the different as my numerator

OpenStudy (anaise):

I don't use the other way.

OpenStudy (anaise):

If that is olay

OpenStudy (anonymous):

its olay

OpenStudy (anonymous):

XD XD

OpenStudy (anaise):

Ken, I have to get a shirt for school tomorrow, adios.

OpenStudy (kendricklamar2014):

Well what I did was: Divide the amount of rose bushes (7) by how many gallons of water she has (5). Which gave me \[\frac{ 7 }{ 5 }\]

OpenStudy (anonymous):

oh well I will be sure to do that thx I'll fan

OpenStudy (anaise):

Srry

OpenStudy (wendy.ivette11714):

Ken is right. Look at this: http://openstudy.com/updates/5540237be4b0a876f245f615